OpenStudy (anonymous):

Find three consecutive even integers such that three times the sum of the first and third is 24 greater than 4 times the second.

OpenStudy (anonymous):

You want to find \(n, (n+1), (n+2)\) such that \(3(n+(n+2)) = 24+ 4(n+1)\). Can you solve this for n?

OpenStudy (anonymous):

no... lol I'm not that good at math

OpenStudy (freckles):

try multiplying a bit

OpenStudy (freckles):

and adding like terms

OpenStudy (freckles):

for example n+(n+2) is the same as n+n+2 which is the same as?

OpenStudy (anonymous):

2n + 2?

OpenStudy (freckles):

yes not do the multiplying part there is lots of places where you can use distributive property that is if you have a(b+c) then this is equal to a*b+a*c

OpenStudy (freckles):

\[3(2n+2)=24+4(n+1)\]

OpenStudy (freckles):

now*

OpenStudy (anonymous):

now what? lol

OpenStudy (freckles):

so you have done the distributing part? and you are asking for what step next? if so please show what you have...

OpenStudy (anonymous):

6n + 6 = 24 + 4n + 4

OpenStudy (freckles):

great now collect like terms because you have some on the right hand side

OpenStudy (anonymous):

6n + 6 = 28 + 4n

OpenStudy (freckles):

right now get your n terms on one side and your terms without n on the opposing side

OpenStudy (freckles):

do this by adding/subtracting terms on both sides

OpenStudy (anonymous):

2n = 22 ?

OpenStudy (freckles):

6n-4n=28-6 2n=22 is right

OpenStudy (anonymous):

n = 11 right?

OpenStudy (freckles):

yes

OpenStudy (freckles):

and oops I just realized the problem said even

OpenStudy (freckles):

the word even is very important here

OpenStudy (freckles):

\[n,n+2,n+4\]

OpenStudy (anonymous):

Oops, my bad.

OpenStudy (freckles):

\[3(n+[n+4])=24+4(n+2)\] solve for n all over again lol

OpenStudy (freckles):

its fine those tiny words get looked over sometimes I should have checked

OpenStudy (anonymous):

would it be 6n + 4 = 24 + 4n + 8?

OpenStudy (freckles):

no 3(2n+4)=6n+12 not 6n+4

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

6n + 12 = 24 + 4n + 8

OpenStudy (freckles):

yes

OpenStudy (anonymous):

6n + 12 = 32 + 4n

OpenStudy (freckles):

great too :)

OpenStudy (freckles):

you are almost there

OpenStudy (anonymous):

2n = 20

OpenStudy (anonymous):

n = 10?

OpenStudy (freckles):

right n=10 n+2=10+2=12 n+4=10+4=14 so the three numbers in question are 10,12,14

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

I don't think @AlixxKK would benefit from a Mathematica script; it isn't really suitable for this level.